### v1

Abstract
We explore a capability of evolution strategies to train an agent with its policy based on a transformer architecture in a reinforcement learning setting. We performed experiments using OpenAI’s highly parallelizable evolution strategy to train Decision Transformer in Humanoid locomotion environment and in the environment of Atari games, testing the ability of this black-box optimization technique to train even such relatively large and complicated models (compared to those previously tested in the literature). We also proposed a method to aid the training by first pretraining the model before using the OpenAI-ES to train it further, and tested its effectiveness. The examined evolution strategy proved to be, in general, capable of achieving strong results and managed to obtain high-performing agents. Therefore, the pretraining was shown to be unnecessary; yet still, it helped us observe and formulate several further insights.
Index Terms:
Evolution strategies, Transformers, Policy optimization, Reinforcement learning
I Introduction
The problem of reinforcement learning is considered one of the most difficult in the field of machine learning. There are many approaches to solving it [1]. Some are based on computing a gradient to optimize the objective, some are derivative-free. One such class of general derivative-free optimization algorithms are evolutionary algorithms [2]. Their subclass of evolution strategies [3] has been proved to be a viable alternative to gradient approaches for the (deep) reinforcement learning [4]. Although the gradient approaches generally have better sample utilization, the evolution strategies are greatly parallelizable. Moreover, evolution strategies have better exploration of possible solutions and the agents trained using these methods are usually more diverse than those trained by the gradient-based algorithms. They can even incorporate techniques that vastly improve the exploration even more, such as searching for novelty instead of (novelty search [5, 6]), or in addition to (quality-diversity [7]) just seeking a better performance. A fairly simple, yet highly effective example of such algorithms for reinforcement learning is OpenAI-ES [8].
On a different note, transformer architecture [9] is lately the go-to solution in the field of neural networks and supervised learning for an ever-growing range of problems. And recently, there have been attempts to reformulate reinforcement learning as a sequence modeling problem and to leverage the capabilities of the transformers in such tasks to obtain a new approach for solving this class of problems, yielding models such as Decision Transformer [10] or Trajectory Transformer [11]. The Decision Transformer was originally introduced as a model for offline reinforcement learning using a supervised learning of sequence prediction, but the authors claim it would function well even in the classical reinforcement learning tasks.
We decided to subject the combination of the evolution strategies and the Decision Transformers to experiments, and test the ability of derivative-free algorithms to train this more complicated and bigger – compared to the simple feed-forward models that had been experimented with in the literature so far – transformer architecture. It might be the case, that for large and complicated models it may be hard, or even almost impossible, to be trained from scratch using an evolutionary approach. Therefore, we wanted to experiment with first pretraining the agent using a supervised learning of sequence prediction on data generated by a smaller, potentially weaker model, which can be trained using any arbitrary reinforcement learning method.
In the following section, we present the background for our experiments. In Section III, we introduce our experiments and show their results, whereas in Section IV, we discuss those results. We then conclude the paper in the last section.
II Background
II-A Evolution Strategies
Evolutionary algorithms are a large and quite a successful family of black-box derivative-free optimization algorithms. They consist of various metaheuristics inspired by nature. One subclass of such nature-inspired optimization methods are evolution strategies. Introduced as a tool for dealing with high-dimensional continuous-valued domains [3], evolution strategies work with a population of individuals (real-valued vectors). In each generation (iteration / population in the iteration), they derive a new set of individuals by somehow mutating (pertrubating) the original population; the new set is then evaluated (with respect to a given objective function), and a new generation is formed based on these new evaluated individuals taking into account their fitness (objective function value).
In order to use an evolution strategy as a reinforcement learning algorithm, we use agents represented by a neural network, or, more specifically, by a real-valued vector of the network weights, as the individuals for the algorithm. Then the only thing needed is to set the fitness of each individual as the mean return of the individual agent from several episodes. Various evolution strategy algorithms were proposed for this purpose [4].
This approach to reinforcement learning has its disadvantages. For example, in order to compute an individual’s fitness, we have to run the entire episodes. Another ailment of this approach is that its sample utilization is low compared to gradient-based methods. That is, in essence, we are capable of extracting more information from a timestep or an episode (a sample) using gradients.
However, evolution strategies also bring a lot of advantages with them. Many of the evolution strategy algorithms are easily parallelizable, and much of the research in the area has been focused on this aspect, yielding us algorithms with a linear improvement of performance when more computing power is used [8]. And because evolution strategies are derivative-free algorithms, we can optimize not just classical smooth neural networks, but our models can even contain some discrete subfunctions or be otherwise non-differentiable.
In this paper, we will be working with OpenAI-ES [8], which is a representative of Natural evolution strategies [12]. The population is represented by a distribution over the agent’s (neural network’s) parameters, the distribution being a Gaussian, whose mean value is the candidate solution to the given problem and from which the offsprings are sampled each generation. These are then evaluated, and their evaluation is used to update the parameters (in our case only the mean value) of the distribution so that we obtain a higher expected fitness for the future samples. This update is performed using an approximation of natural gradient (whence the name Natural evolution strategies). In our case, this is obtained by renormalizing (rescaling) the update with respect to uncertainty. (In general, we would have to compute an inverse of a so-called Fisher information matrix, which would then be used to multiply the gradient estimate by. However, as demonstrated, e.g., in a paper by Pierrot, Perrin-Gilbert and Sigaud [13], in our case, when we derive the parameter updates from the Gaussian distribution with the same variance for each parameter, dividing by the variance (the rescaling w.r.t. uncertainty) achieves approximately the same result.) The algorithm is also designed in such a way that it is highly parallelizable with interprocess communication kept at bare minimum. For more details or for a discussion of the design choices, we refer our readers to the original paper [8].
II-B Transformers
Transformers have surged as the state-of-the-art neural architecture for numerous tasks of supervised learning in recent years. It all started with their introduction as a new sequence-to-sequence architecture [9], as an alternative to recurrent neural networks. Since then they have yielded great results in natural language processing (NLP) tasks, fueling even the current surge of chatbots, and they even got adapted, e.g., to image recognition [14]. As a rule of thumb, they seem to have a great generalization power; the greater the larger the model employed. However, they also require a lot of training data to achieve these great results.
The transformers interlace classical feed-forward layers with self-attention layers. (In practice, there are also residual connections and layer normalizations to stabilize the training.) And it is the self-attention to which the transformers owe their success. For each element of an input sequence, the self-attention constructs a "key", a "query", and a "value" using fully connected layers of neurons. Then, to compute an $i$-th element of an output sequence on the same position, it takes a combination of all the values, each weighted proportionally to the product of the query belonging to the given ($i$-th) position and a key corresponding to the value – hence combining the information from the whole input sequence to produce every single element of the output sequence, as shown by the following equation.
$output_{i}=\sum_{j=1}^{n}\text{softmax}\left(query_{i}^{T}\cdot all\_keys\right)_{j}\cdot value_{j}$
We can, of course, use a mask and for every element hide the part of the sequence that is behind the element, and so use just the information that came before in the sequence to derive the output element. This is called a causal masking, and such a transformer that uses this masking we then call a causal transformer.
In the problem of reinforcement learning, we want the agent to choose actions at individual timesteps, such that it maximizes its return. This can be, however, viewed as a sequence modeling problem. And then, naturally, the best sequence processing architecture available, the transformer, comes into focus. Thus, the Decision Transformer was introduced [10].
Its main idea is that we want the agent’s policy to produce an action based on not just the last observation, but the whole history (or the part which squeezes into the context window) of past observations and undertaken actions. And to have some way to affect the agent’s performance, we add a conditioning on a return-to-go, which is a return we want to obtain from a given step until the end of the episode.
So, let us take a look at the proposed architecture itself. The Decision Transformer consists of a causal transformer; embeddings for returns-to-go, observations (states of the environment), and actions; position encoder; and a linear decoder to transform the output of the transformer into actions, as shown in Figure 1.
Figure 1: Decision Transformer architecture [10]
Every timestep, we feed the model with a sequence of past triplets return-to-go, observation, and action performed, adding the current return-to-go and observation (and a placeholder for the yet unperformed action). They get embedded by their respective embeddings, the positional encoding is added, and everything is then fed through the transformer. We then decode the last output state of the transformer to obtain the action to be carried out. An important difference from the usual transformer is that every part of the timestep triplet, meaning return-to-go, observation, and action belonging to one particular timestep, shares one positional encoding, as opposed to the classical transformer, where every input sequence element gets its own. Returns-to-go are constructed in what could be called a recursive manner. The user has to supply the original one for the first timestep (which has a meaning as the desired performance of the model, in other words target return), while for all the following timesteps the return-to-go is constructed by subtracting the last reward obtained from the return-to-go belonging to the previous timestep.
III Experiments
From what we saw in the previous section, the evolution strategies belong to the state-of-the-art algorithms for solving the reinforcement learning problem, whereas the transformers embody a new possibility for how to approach this problem, as well as a potentially strong architecture of the neural network representing the agent’s policy. Therefore, it is just a logical next step to try and combine these two concepts, to test the capability of the transformers to be trained by evolution strategies in the setting of reinforcement learning.
We provided our own implementation of the OpenAI-ES  [^1] Our code, together with the data gathered during all the experiments can be found at the following address: https://github.com/Mafi412/Evolution-Strategies-and-Decision-Transformers and began by a replication experiment of the original paper, which also serves as a correctness check for our code. Here, we tested the algorithm on a classical feed-forward network in the MuJoCo [15] Humanoid environment using OpenAI Gym [16], whereupon we followed up by proceeding to test the performance of the evolution strategy with the Decision Transformer in the same environment. We then examined an idea of pretraining the transformer by a behavior cloning of some smaller and easily trained – yet possibly weaker – model before training the transformer by the evolution strategy. The last experiments we present are studying the particulars of training a larger model in a different environment with a higher-dimensional image input, which is Hero, one of the Atari games [17].
In all the experiments with Decision Transformers, we used the same hyperparameter values for the model as used in the original paper [10] for the respective environments. For comparison, the model sizes are 166 144 parameters for the feed-forward model from the original OpenAI-ES paper [8], 825 098 parameters for the Decision Transformer for MuJoCo Humanoid environment and 2 486 272 parameters for Atari games environment.
For all experiments, we conducted ten runs of the training and aggregated the results in our plots. For every run of each of the experiments, 300 workers were used utilizing 301 CPU cores (with one being a master handling synchronization, evaluation, and saving the agent). The desired returns passed to all the Decision Transformer models at the beginning of each episode were 7000 for the Humanoid Decision Transformer and 8000 for the Atari Decision Transformer.
Unless stated otherwise in the individual experiment descriptions, all the other hyperparameter values can be found as default values in our codebase.
(a) Evaluation results
(b) Population fitnesses
Figure 2: OpenAI-ES used on a simple feed-forward model for a MuJoCo Humanoid simulation in ten runs of the experiment. In Figure 2(a), we can see the results of evaluations of the current solution model after every iteration of the algorithm. Figure 2(b) shows the fitness values of the whole populations throughout the iterations (mean and percentile interval with standard width, ranging from 2.5 to 97.5 percentiles).
Our Implementation
Mainly due to different computational resources available to us – we needed to use the MPI framework instead of EC2 (AWS) – as well as the need to interchange the models being trained and the environments for the evaluation more easily and with a greater degree of freedom, we decided to create our own implementation of OpenAI-ES. Yet, in the process of creating this implementation of ours we noticed some inconsistencies between the original paper introducing the algorithm [8] and their provided code. We decided to adhere more to the paper when these dissimilarities occured and here we state our reasoning for the two most significant differences.
First and foremost, in the paper [8], the use of a weight decay is mentioned as a form of keeping the effect of newly added updates still significant enough. After every update, we decay the weights of our model, which is the distribution mean. Therefore, it would not happen that there could be weights grown to such sizes that a newly added noise practically does not change the function of the model. This reasoning tells us why we want to use preciselly the weight decay. Nevertheless, in the original implementation L2-regularization is used instead. This may be all right for some optimizers, as for, e.g., SGD or SGD with momentum it is virtually the same as the weight decay (just rescaled by the learning rate). However, as mentioned in paper Decoupled Weight Decay Regularization [18], it is not the same, for example, for ADAM optimizer, which is nonetheless exactly the one used in the original experiments. Our implementation, on the other hand, remains true to the original paper and uses proper weight decay. Nevertheless, we employ SGD with momentum in our experiments, as it demonstrated superior efficacy in our setting during our experiments.
Secondly, during the update of the distribution parameters, a gradient estimate should be normalized with respect to the uncertainty using a division by the standard deviation, so that it becomes a natural gradient estimate. Despite that, in the original OpenAI code, this is not done. There, this is probably hidden in the value of a learning rate (which, as a result, differs from ours). However, then the learning rate and the noise deviation are unnecessarily coupled hyperparameters.
In addition to the implementation of OpenAI-ES, we made a few slight modifications even in the Decision Transformer code. The most significant one concerns keeping data from the past to be used in future inputs. In the original code, the authors do not clip sequences already longer than the context window; those are hence allowed to grow unconstrained in the memory. The new data are always just appended to the list, keeping the whole history, and the whole list is passed to an inference function, which only then clips the sequence to the required length (without modifying the data stored) and passes it to the transformer. This proved to be no problem in MuJoCo environment, but for some reason leads to a severe slowdown in the Atari environment. Thus, we keep the stored sequence cropped to match the context length in the Atari version of Decision Transformer.
(a) Evaluation results
(b) Evaluation results with half the population
Figure 3: OpenAI-ES used on a Decision Transformer model for a MuJoCo Humanoid simulation in ten runs of the experiment. Figure 3(a) shows us the evaluation results of the current solution in each generation of the evolution strategy, whereas in Figure 3(b) we can see the evaluations for the experiment using just half the population.
(a) Evaluation results
(b) Evaluation results with half the population
Figure 4: OpenAI-ES used on a pretrained Decision Transformer model for a MuJoCo Humanoid simulation in ten runs of the experiment. In Figure 4(a), the evaluation results of the current solution of each iteration are shown, and Figure 4(b) presents the evaluation results for this experiment using only half the population.
III-A Feed-Forward - Humanoid
As stated before, the first experiment serves as a replication experiment for the original OpenAI-ES paper [8], as well as a sanity check of our implementation. It consists of training a feed-forward model using OpenAI-ES in Humanoid environment. We can see in Figure 2 that our implementation is functional and is comparable in performance with the original implementation. The progression of the training shown in Figure 2(a) also serves as a benchmark against which the results of other experiments can be compared.
III-B Decision Transformer - Humanoid
In this experiment, we inspected the ability of the given evolution strategy to train even the more complicated model, which the Decision Transformers are. We trained a Decision Transformer model using OpenAI-ES, again in the Humanoid environment. Because the model is almost five times larger, we quadrupled the size of the population the algorithm works with. The results are shown in Figure 3(a).
We also explored whether it is necessary to use such a substantial population or whether half of it would be sufficient. The results of this ablation study can be seen in Figure 3(b). We can see that even though the algorithm is sometimes still capable of training the model, more often than not it does not succeed in doing so.
(a) Evaluation results
(b) Evaluation wall-clock times
Figure 5: OpenAI-ES used on a Decision Transformer model for Atari game Hero in ten runs of the experiment. Figure 5(a) shows us the evaluation results after every iteration of the algorithm. In Figure 5(b), we can see the wall-clock time the algorithm needed each iteration in seconds.
III-C Decision Transformer - Humanoid - Pretrained
The core idea behind this experiment is that if we have a really large model with many parameters, it might be quite hard for the algorithm to gain enough information from just sampling in a neighborhood of a randomly initialized model. Hence what we might try to do is first pretrain the model using behavior cloning towards some smaller, easily trained yet possibly weaker model and only then utilizing the evolution strategy to further improve the large pretrained model.
Yet, this approach has its costs. First, the OpenAI-ES uses virtual batch normalization (VBN) [19]. This changes the inputs to the model as the training progresses. So, either we would have to use VBN even during the pretraining, or we cannot use the VBN during the following training with the evolution strategy. We opted for the second possibility, as it is more general and allows us to showcase a further training of any model we might have using the evolution strategy, even though the first option would be more favorable for the algorithm, as the VBN improves its reliability [8] – it should, however, be most important, when the model is mostly random at the beginning of the training, according to the original authors.
Second, using the same values for learning rate and noise deviation hyperparameters as when training from scratch leads to complete randomization of the model at the beginning of the training and then the training proceeds as if from scratch. So, for the pretraining to have an effect, we have to decrease both values (to 0.01 in our experiment). This, in turn, leads to a slower progression of the training.
Hence, again in the Humanoid environment, we pretrained the Decision Transformer using a behavior cloning to a feed-forward model trained using SAC algorithm [20], before training it further using OpenAI-ES. The feed-forward model and the pretrained Decision Transformer had a fitness value around 4. The results of the subsequent training of the pretrained Decision Transformer using the evolution strategy are shown in Figure 4(a).
Again, we explored the possibility of utilizing a smaller population for this training. As can be seen in Figure 4(b), this time, although the reliability of the training is again diminished, most of the time the algorithm works surprisingly well compared to how it works with the full population.
Next, we tried using standard (i.e. the same as in the previous experiment without the pretraining) values for learning rate and noise deviation hyperparameters both with and without the VBN. Without the VBN the training showed no progression at all, and with the VBN the training started to show similar patterns as the one from the previous experiment, yet still was maybe a bit slower. The corresponding plots are quite uninteresting, but together with the underlying experiment data may be found along the code and other data for this paper.
III-D Decision Transformer - Atari
As a final experiment, we tested the ability of OpenAI-ES to train an even larger Decision Transformer in Atari games environment, more specifically, in the game Hero. Again, since the model is approximately three times bigger than in the previous experiments, we further doubled the population size for the algorithm. Even though Figure 5(a) shows that a gradual improvement of the model really does occur, Figure 5(b) reveals that it is accompanied by quite a significant increase in the time required to process a single iteration. This is due to slower model inference, as the model is substantially larger, as well as longer episodes in the environment.
In Figure 5(b), we can see an interesting pattern with a spike at the beginning of the training, so let us just note that it is something domain dependent. Here is our attempt to explain what is most likely going on. At the beginning, the agent performs random actions, which leads it to repeatedly blow itself up, hence we get relatively short episodes. Then it learns that planting the bombs randomly is not a good idea and stops ending the episodes prematurely. And then it gradually learns how to get further and further through the levels of the game, but this comes at the cost of it occasionally blowing itself up again, yet this time it at least serves its progression through the game.
IV Discussion
As we can see by comparing Figures 2(a) and 3(a), the larger Decision Transformer model has an fitness advantage during the training over the smaller feed-forward model. But what is more important for this article, OpenAI-ES appears capable of training it without substantial problems, although the process takes longer. However, the training is less robust, as evidenced in Figure 3(a), which shows that during one experiment run, the model failed to achieve any significant improvement. For greater robustness, we would probably have to further increase the population size.
Even though the parallelization of the algorithm significantly improves the time required by the algorithm to train the model, on such a massive scale (or even more in some potential use cases) it still places substantial demands on the computing hardware, or rather on the reliability of communication between processes.
Regarding our specific usage of the algorithm on Decision Transformers, we need to settle one thing. As we introduced in Section II-B, Decision Transformer operates with return-to-go tokens, which in the original paper [10] serve as a desired return the agent should achieve in a remainder of an episode. This works well when training in the original offline reinforcement learning manner, but when using the evolution strategy to train the Decision Transformers in our online setting, there is no pressure to take these tokens into account, and thus the tokens are then ignored. Nonetheless, we believe that only a minor modification of the algorithm may induce the agent to pay attention to these tokens. Let us propose an outline of such modifications for possible future work. Each evaluation would consist of several subevaluations; each subevaluation would be assigned a desired return sampled from a $\mathcal{N}(\mu,\sigma)$, a normal distribution centered at $\mu$ with variance $\sigma$. The value of $\mu$ should be computed from the returns obtained during the previous iteration in such a manner that it is within an interval between the best return obtained and the mean return obtained during the last iteration, so an improvement is gradually made. The variance $\sigma$ would be a hyperparameter, or it could possibly be a variation of the returns obtained during the last iteration. The fitness of an individual would not be the mean of its returns obtained, as is the case in OpenAI-ES, but rather a decreasing function of the absolute values of the differences between the return obtained and the desired return in each subevaluation; ergo, the larger the differences, the smaller the fitness. This might be, e.g., the negative of a weighted sum of the differences.
Now, let us examine the experiments that utilize the pretraining described in Section III-C. As we noted in the section when explaining the hyperparameter values, the pretraining does not perform very well. In order for us to be able to use the pretrained model as a base for further training – even if we would be using VBN already during the pretraining – we have to substantially reduce the ability of the algorithm to further improve the model, because we have to change the learning rate and noise deviation hyperparameters to smaller values. Hence, we get slower learning and less exploration per iteration.
Even when doing so, we can see in Figure 4 that the model first gets "broken" – its performance worsens during first few iterations – before it starts to improve again. Nevertheless, the improvement comes in terms of iterations much sooner than when no pretraing is utilized. Furthermore, we can observe by inspecting rollouts of both the initial and final models that those two show similar gaits. This similarity gets even more accentuated, when compared to the diversity in final behaviors that are yielded from the experiment in Section III-B, so when the algorithm is run without the pretraining. Therefore, we can clearly see that when initialized by the pretrained model the training clearly builds upon the seeded model, despite its initial deterioration.
A reason for the above-mentioned deterioration at the beginning of the training from pretrained model, as well as the need for the aforementioned change of hyperparameters is, we believe, the following: Gradient training – of which the behavior cloning is a representant – follows a different overall goal than the evolution strategy. It strives for the best possible parameters of the model it trains with regard to its loss function. In our case of behavior cloning, this means that we want to get our model to behave exactly as the one we are cloning towards. Nonetheless, this tells us nothing about how the model behaves when we perform a slight perturbation of parameters. Hence, after the perturbation, the model is likely to no longer function properly – or at least as well – anymore. So we can say that the model is brittle in a sense that even a slight change of parameters is likely to somehow break the model. This is, however, in stark contrast to how the evolution strategy operates and what it tries to achieve. The goal of the evolution strategy is not only that its current model performs as well as possible, but that we get the best possible behavior in expectation when sampling model parameters from its current distribution. Thus, the models trained by the evolution strategy might be considered more robust – that might help, e.g., in areas where we train on precise hardware, but the model needs to be afterwards deployed on a cheaper hardware with less precision, where the weights of the model need to be rounded. Ergo, the algorithm needs to first, let us say, "robustify" the model and only then it can train it towards a better performance.
Regardless, in the end, the results shown in Section III-B suggest that the pretraining is not really necessary. But then what explains why the problem stated at the beginning of Section III-C does not occur in practice? In Figure 2(b), we can see that the algorithm improves not only by shifting model weights towards the few better performing individuals of the population, but mostly by shifting the weights away from the worst performing individuals, as the majority of the population consists of worse individuals. This might indicate why the algorithm does not really struggle with training larger and more complex models, where it is even harder to stumble onto a working solution just by adding some noise to the model parameters.
Finally, let us just briefly comment on Section III-D. Yes, the OpenAI-ES is capable of training even such larger models, but it just takes too long. Of course, this could be easily solved by even more massive paralelization over even more processes and CPUs – although 300 workers is already quite a lot and we believe there is a limit on the scaling. But the question lingers: How fast would a gradient algorithm – even more so some Temporal-Difference based [21] (Actor-Critic like [22]) – be able to train these large models? And on top of that, gradient algorithms work best on GPUs, which one might also use to accelerate an inference in such large models. Meanwhile, evolution strategies work with a lot of CPUs instead, which are mostly better for environment simulations, but worse for the model inference, therefore we have a slower inference during rollouts – the slower the larger the model we have. So, in essence, where is the boundary, when the length of an episode in an environment and the size of a model favor gradients over evolution? (And yes, this question is – just as similar questions have always been in history of artificial intelligence and even more so of machine learning – deeply dependent on a current state of hardware available. But we believe the hardware cannot be improved indefinitely.) An indirect answer to this question might be utilization of hybrid algorithms that combine strengths of both approaches, the gradient algorithms, and the evolution [23].
V Conclusion
We have investigated the ability of OpenAI-ES (and, in extension, of similar evolution strategies) to train larger and more complex models than the standard feed-forward ones used in the literature so far in reinforcement learning. We have also suggested a method to help this training by pretraining the model using behavior cloning towards some smaller, possibly weaker, yet easier to train model. However, the evolution strategy proved to be capable of training large models, whereas the pretraining showed multiple flaws during the experiments. Still, it helped us shed light onto the field of hybridizing gradients and evolution strategies, as we have discussed based on our experiments. It helped us understand why the sequential combination of first training using gradients and then proceeding to train using similar distribution-based evolution strategies will not work well. And it also served to support the claim that the distributional evolution strategies like OpenAI-ES yield robust models, which are less prone, e.g., to rounding the weights due to low-precision hardware they might be deployed on.